OpenStudy (anonymous):

Find the volume of the sphere to the nearest whole number. Use p = 3.14. A. 131 in.3 B. 393 in.3 C. 4,187 in.3 D. 523 in.3

OpenStudy (anonymous):



OpenStudy (anonymous):

help plzzzzzzz

jimthompson5910 (jim_thompson5910):

V = (4/3)*pi*r^3 V = (4/3)*3.14*(5)^3 V = ???

OpenStudy (anonymous):

523.333333=523 is that right

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thank u

jimthompson5910 (jim_thompson5910):

np